# Frozen Transformers in Language Models Are Effective Visual Encoder Layers

### v1

@@ 4.4 Motion Forecasting
Task statement.   We select motion forecasting as an example of a non-semantic task. It is safety-critical for autonomous driving and capitalizes on the understanding of dynamics, agent-agent interaction, and agent-lane relationship. The input commonly includes the historical trajectories of agents and way-points of lane segments, which are both represented in polylines on the bird’s-eye view (BEV). The desired output is a set of $K$ most possible future trajectories, where the multi-future setup explicitly accounts for the uncertainty of urban traffic. We conduct experiments on Argoverse (Chang et al., 2019). The evaluation metrics are minimum average displacement (ADE), minimum final displacement (FDE), and miss rate (MR), which calculate the errors of predictions from different aspects and are better at lower values.
@@
Design.   We apply the frozen LLM transformer to VectorNet (Gao et al., 2020) and mmTransformer (Liu et al., 2021). Both of the models first convert the polylines of agents and lanes into token embeddings and their encoders leverage a series of attention blocks to process these agent and lane tokens. Our LLaMA transformer block is situated behind their encoder and processes these agent and lane tokens before feeding the feature of the target agent into the decoder. Demonstration and details are in Sec. C.4.
Analysis.   According to Table 4, the models with LLaMA forecast better trajectories, which is reflected by smaller distances to the ground truth. However, we notice that the improvement is less significant compared with semantic tasks, which reflects the preference of LLM transformers for encoding rich semantics over abstract object movements.
@@ 4.5 Vision-Language Tasks
4.5.1 2D Vision-Language tasks
Task statement.   The benefits of frozen LLM transformers for visual encoding are not limited to purely visual tasks. We experiment with 2D vision-language (VL) tasks, including visual question answering (VQA) on VQAv2 (Goyal et al., 2017c) and zero-shot image retrieval on Flickr30k (Plummer et al., 2015). In both of the tasks, the input to the model includes both images and texts. Therefore, they evaluate the ability of the frozen language transformer to fuse the multi-modal vision-language features for downstream tasks.
Design.   We adopt the widely-used METER (Dou et al., 2022) as our baseline. It first extracts uni-modal features for images and texts, then fuses cross-modal features, and finally decodes the output from the cross-modal features. Similar to the previous design that situates the LLM transformer block before the decoder, we insert the frozen transformer right between the cross-modal fusion and decoder modules of METER, which is the end of METER’s encoder. An intuitive illustration is in Fig. H. In the experiments, our setup follows Shi et al. (2023): initialize the image encoder with CLIP-B/32 (Radford et al., 2021) and text encoder with RoBERTa (Liu et al., 2019), and then jointly finetune all the modules on VQAv2 or Flickr30k, except for the frozen LLM transformer. More implementation details are covered in Sec. C.5.
Analysis.   As in Table 4(a), both of the 2D VL tasks are significantly enhanced with the LLaMA transformer, including the improved quality of question answering on VQAv2 and retrieval recall on Flickr30k. This evidence supports the potential of using a frozen LLM transformer for encoding multi-modal features.
@@
Model VQAv2 (Test-dev) Flickr30k (Val)
@@
4.5.2 3D Visual Question Answering
Task statement.   We extend our proposed idea into 3D VQA, which requires comprehending an input 3D scene represented by point clouds or multi-view images and then answering questions. 3D VQA challenges the ability to ground language in 3D contexts, where the model needs to understand the three-dimensional visual context in the form of point cloud in synergy with the language context. We conduct our experiments on the SQA3D (Ma et al., 2023) dataset, comparing with baseline methods and state of the arts (Ma et al., 2023; Azuma et al., 2022) on the exact match (EM) metric.
Design.   Similar to 2D VL models, the baseline of SQA3D also adopts the modular design of uni-modal encoding and multi-modal feature fusion before decoding the output. It processes the textual input with LSTM (Hochreiter & Schmidhuber, 1997) and 3D point clouds with VoteNet (Qi et al., 2019), then conducts VL fusion before feeding into the decoder. Here, we add the frozen LLM block after the VL fusion, which is consistent with our 2D VL design (Sec. 4.5.1). More details of experiments are in Sec. C.6.
Analysis.   According to Table 4(b), adding a frozen LLM transformer effectively enhances the ability of the model to fuse point cloud and language features and ground them for 3D question answering. We further provide detailed breakdown metrics in Table B, where adding the LLM transformer benefits most of the scenarios.
@@ 5 Analysis on LLM Transformers for Visual Tasks
We justify our design choices (Sec. 5.1) and illustrate the wide applicability of our framework to various LLMs and transformer layers (Sec. 5.2). Our investigation also discovers that sufficiently large LLMs are the premise of benefiting visual encoding with a frozen transformer (Sec. 5.3).
@@ 5.1 Ablation Study on Design Choices
Table 6: Ablation study on model capacity and finetuning.
Model Capacity.   Regarding the wide applicability of frozen LLM transformers, we question if the improvement mainly comes from the increased capacity of the linear layers $\mathbf{F}_{L}^{1}$ and $\mathbf{F}_{L}^{2}$, instead of the pre-trained weights in LLM transformers $\mathbf{F}_{LM}$. To analyze model capacity, we create ViT-S-MLP, which has identical trainable parameters compared with ViT-S-LLaMA. Concretely, ViT-S-MLP removes the LLM block $\mathbf{F}_{LM}$, and then inserts a GeLU activation (Hendrycks & Gimpel, 2016) and layer normalization (Ba et al., 2016) between $\mathbf{F}_{L}^{1}$ and $\mathbf{F}_{L}^{2}$. It also adopts the identical training procedure as ViT and ViT-LLaMA in Sec. 4.1 for a fair comparison. The results are summarized in Table 6: the ViT-S-MLP has better performance than ViT-S due to its increased capacity, but the improvement is only about half of ViT-S-LLaMA. Therefore, the LLM transformer weights are crucial for the improvement and the observed benefits are not mere consequences of an increased model capacity.
@@
Finetuning.   We further verify whether finetuning the language transformer (ViT-S-LLaMA-FT) is better than freezing it. As in Table 6, finetuning decreases the performance compared with ViT-S-LLaMA. We analyze this phenomenon by visualizing the loss curves in Fig. F and training under an additional setting of 100 epochs. Although finetuning improves performance under short training (100 epochs in Table A), it hurts the accuracy when trained sufficiently: in Fig. F, ViT-S-LLaMA-FT shows lower training loss but relatively larger validation loss, which indicates overfitting. Thus, our observation demonstrates the challenges of training large transformers, and we accordingly freeze the LLM transformers in our design because of its simplicity and effectiveness.
@@ 5.2 Varying LLM Transformer Layers
We discover that LLM transformers influence visual representation learning significantly under our framework, even though they have identical capacities. Specifically, we use transformer blocks from diverse depths of LLaMA-7B (Touvron et al., 2023) and OPT (Zhang et al., 2022) onto ViT-S. The models are trained on the ablation study setting of 100 epochs. (More details in Sec. C.7.) As shown in Fig. 2, the type of layer significantly changes the performance. These experiments also validate that our framework is applicable to various LLMs and transformer layers, and highlight the importance of selecting proper transformer layers. We additionally observe that the last LLM layers consistently improve the performance although they might not be optimal.
5.3 Varying LLM Transformer Scales
Model Acc
ViT-S 75.25
+ OPT-125M 71.63
+ OPT-350M 71.56
+ OPT-1.3B 75.62
+ OPT-2.7B 75.74
+ OPT-6.7B 76.29
Table 7: Accuracy improves with larger transformer.
[…]
As indicated by the results in Table 7, the benefits of frozen language transformer grow with an increasing capacity of OPT transformers. The added transformers enhance the performance only with sufficient sizes (1.3B, 2.7B, 6.7B) and hurt the accuracy when the sizes are small (125M, 350M). Therefore, the phenomena of LLM transformers enhancing visual tasks only “emerges” at sufficient scales.
@@ 6 Information Filtering Hypothesis
This section aims to explain how a frozen and pre-trained LLM transformer benefits visual tasks. Intuitively, our hypothesis can be stated as:
Information filtering hypothesis.   The pretrained LLM transformer functions as a “filter” that distinguishes the informative tokens and amplifies their contribution for the prediction, in the form of enlarged magnitudes or frequencies in the feature activation.
We first derive this hypothesis in the context of image classification (Sec. 6.1), then provide quantitative investigation (Sec. 6.2) and discuss the observation on other tasks (Sec. 6.3). Finally, we discuss the connection with usable information and limitations are discussed in Sec. 6.4 and we include more details in Sec. A.
@@ 6.1 Qualitative Derivation of Information Filtering Hypothesis
As clearly demonstrated in Fig. 2(a), the token activation better captures the regions of target objects after adding the LLM transformer, especially the magnitudes of $\mathbf{F}_{L}^{2}$ and frequencies of $\mathbf{F}_{LM}^{A}$. Their tendency of segmentation is a surprising discovery because emergent segmentation only exists in self-supervised learned (Caron et al., 2021) or specially-designed (Shi et al., 2023) ViTs. More importantly, the activation’s concentration on the target object directly supports our hypothesis as evidence of selecting the informative tokens.
@@
Noisy attention scores.   In contrast to the feature activation, attention scores struggle to capture the relevant visual tokens for prediction. We investigate the attention scores between the $\mathtt{CLS}$ token and visual tokens in the last transformer block, which are the last self-attention block in $\mathbf{F}_{E}$ for ViT and the transformer of $\mathbf{F}_{LM}$ for ViT-LLaMA, respectively. Ideal attention scores that distinguish the target object should exhibit object segmentation patterns like DINO (Caron et al., 2021). However, supervised ViT models commonly have noisy attention scores (left part in Fig. 2(b)). Although ViT-LLaMA illustrates the ability of emergent segmentation in a few attention heads, most of the attention scores also suffer from scattering and noisiness. These observations contrast the feature activation and indicate that the benefits of LLM transformers cannot be simply attributed to attention scores, since attention scores fail to reliably contribute correct visual tokens.
Deriving the amplification of informative tokens.   According to our visualization in Fig. 2(a), the frozen LLM transformer distinguishes the informative tokens. Intuitively, such tokens naturally benefit the downstream decoding, but this is straightforward only when the decoder directly takes the visual tokens as input. As a counterexample, ViT utilizes a $\mathtt{CLS}$ token for classification, and the visual tokens output by $\mathbf{F}_{L}^{2}$ are not the input to the decoder and always receive zero gradients during training. To bridge the gap in $\mathtt{CLS}$ token scenarios, the second half of our hypothesis is necessary: the frozen LLM transformer amplifies the contribution of informative tokens.
Concretely, the calculation of $\mathtt{CLS}$ token is,
@@
where $V$ denotes visual tokens, $w_{v}$ denotes weight of visual token $v$. Eqn. 3 describes the process of (1) aggregating visual tokens $z_{L}^{1}[v]$ guided by the attention scores $w_{v}$; and (2) sequentially flowing through the remaining layers, including the MLPs in LLM transformer’s attention head $\mathbf{F}_{LM}^{A}$, feedforward network $\mathbf{F}_{LM}^{F}$, and the second linear layer $\mathbf{F}_{L}^{2}$. To concentrate on explaining visual tokens, Eqn. 3 also slightly simplifies self-attention by removing $\mathtt{CLS}$ token from the right-hand side.
In Eqn. 3, the useful visual tokens are not reliably attributed to $\mathtt{CLS}$ tokens because the attention scores $w_{v}$ are observed to be noisy. Therefore, our objective is to bridge the visual tokens in Eqn. 4 below, which are informative (as in Fig. 2(a)), to the final feature representation $z_{L}^{2}[\mathtt{CLS}]$.
$z_{L}^{2}[v]=\mathbf{F}_{L}^{2}\cdot\mathbf{F}_{LM}^{F}\cdot\mathbf{F}_{LM}^{A}(z_{L}^{1}[v]),\ \text{where}\ v\in V.$ (4)
This pursuit leads to our hypothesis that the frozen LLM blocks amplify the fraction of informative token features in addition to distinguishing them. We express the hypothesis below in a formal way, which is a simple change to Eqn. 3:
@@
Eqn. 5 holds equal under the special case of $\mathbf{F}_{L}^{2}\cdot\mathbf{F}_{LM}^{F}\cdot\mathbf{F}_{LM}^{A}$ being linear transformation. This equation explains how the informative tokens in Fig. 2(a) are implicitly reflected and supervised in the $\mathtt{CLS}$ token. It further clarifies how the informative tokens can benefit the final prediction though the attention scores $w_{i}$ are noisy.
As a brief remark, our derivation builds upon two observed pieces of evidence: (1) visual tokens concentrating on informative regions; and (2) noisy attention scores. These lead to the first half of our hypothesis and explain the benefits when visual tokens are direct input to decoders. By further bridging visual tokens to $\mathtt{CLS}$ tokens, we propose that the inserted LLM transformer amplifies the effects of informative tokens. A more thorough version of the derivation is in Sec. A.1.
@@ 6.2 Quantitative Evidence
Figure 4: Pseudo masks from ViT-LLaMA’s features ($\mathbf{F}_{L}^{2}$) have larger mIoU than attention scores and ViT, which supports our hypothesis that the frozen transformer distinguishes informative regions.
@@
Finally, we summarize the mIoU statistics for feature activation and attention scores in Fig. 4. As demonstrated, both ViT-S-LLaMA and ViT-B-LLaMA have better mIoU of pseudo-masks than attention scores. This directly supports our hypothesis that the features $\{z_{L}^{2}[v])|v\in V\}$ contribute more reliably than attention scores $\{w_{v}|v\in V\}$. We additionally notice that the pseudo-masks from ViT-LLaMA generally have larger mIoU compared with ViT, which reflects the benefits of training ViTs with a frozen LLM transformer. The advantage of the feature in the first linear layer $\mathbf{F}_{L}^{1}$ also reveals that training with our framework is beneficial to even earlier stages of features. However, we would like to clarify that the pseudo-masks from either magnitude or frequency activation are intuitive but lossy measures to quantify feature quality because neural networks can encode information into other formats. Therefore, developing better measurements to analyze individual network layers will be meaningful for future work.

### v2

@@ 4.4 Motion Forecasting
We select motion forecasting as an example of a non-semantic task. It is safety-critical for autonomous driving and capitalizes on the understanding of dynamics, agent-agent interaction, and agent-lane relationship. The input commonly includes the historical trajectories of agents and way-points of lane segments, which are both represented in polylines on the bird’s-eye view (BEV). The desired output is a set of $K$ most possible future trajectories. We conduct experiments on Argoverse (Chang et al., 2019). The evaluation metrics are minimum average displacement (ADE), minimum final displacement (FDE), and miss rate (MR), which calculate the errors of predictions from different aspects and are better at lower values. We apply the frozen LLM transformer to VectorNet (Gao et al., 2020) and mmTransformer (Liu et al., 2021). They first convert the agents and lanes into features, and then our LLaMA transformer block processes these agent and lane tokens. Demonstration and details are in Sec. C.4.
According to Table 4, the models with LLaMA forecast better trajectories. However, we notice that the improvement is less significant compared with semantic tasks, which reflects the preference of LLM transformers for encoding rich semantics over object movements.
@@ 4.5 Vision-Language Tasks
2D vision-language tasks.   The benefits of frozen LLM transformers for visual encoding are not limited to purely visual tasks. We experiment with 2D vision-language (VL) tasks, including visual question answering (VQA) on VQAv2 (Goyal et al., 2017c) and zero-shot image retrieval on Flickr30k (Plummer et al., 2015). We adopt the widely-used METER (Dou et al., 2022) as our baseline. It extracts uni-modal features for images and text, fuses cross-modal features, and decodes the output from the cross-modal features. We insert the LLM transformer block after the cross-modal fusion. An intuitive illustration is in Fig. H. During training, our setup follows Shi et al. (2023): initialize the image encoder with CLIP-B/32 (Radford et al., 2021) and the text encoder with RoBERTa (Liu et al., 2019), and then fine-tune on VQAv2 or Flickr30k. Details are in Sec. C.5. As in Table 5(a), both of the 2D VL tasks are significantly enhanced with the LLaMA transformer. This evidence supports the potential of a frozen LLM transformer for multi-modal tasks.
@@
Model VQAv2 (Tes-dev) Flickr30k (Val)
@@
3D visual question answering.   We extend our proposed idea into 3D VQA, which requires comprehending an input 3D scene represented by a point cloud or multi-view images and then answering questions. 3D VQA challenges the ability to ground language in 3D context. We conduct our experiments on the SQA3D (Ma et al., 2023) dataset, comparing with baseline methods and state of the arts (Ma et al., 2023; Azuma et al., 2022) on the exact match (EM) metric. We follow the baseline SQA3D to process the textual input with LSTM (Hochreiter & Schmidhuber, 1997) and 3D point clouds with VoteNet (Qi et al., 2019). Here, we add the LLM block after the VL fusion, which is consistent with our 2D VL design (Sec. 4.5). More details are in Sec. C.6. According to Table 5(b), adding a frozen LLM transformer effectively enhances the QA ability of models. The full comparison with detailed breakdown metrics is in Table E.
@@ 5 Analysis on LLM Transformers for Visual Tasks
We justify our design choices (Sec. 5.1) and illustrate the wide applicability of our framework to various LLMs and transformer layers (Sec. 5.2). Our investigation also discovers that sufficiently large LLMs are the premise of benefiting visual encoding with a frozen transformer (Sec. B.3) and discusses the place to insert LLM blocks (Sec. B.4).
@@ 5.1 Ablation Study on Design Choices
Table 6: Ablation study on model capacity and fine-tuning.
Model capacity.   Regarding the wide applicability of frozen LLM transformers, we question if the improvement mainly comes from the increased capacity of the linear layers $\mathbf{F}_{L}^{1}$ and $\mathbf{F}_{L}^{2}$, instead of the pre-trained weights in LLM transformers $\mathbf{F}_{LM}$. To analyze model capacity, we create ViT-S-MLP, which has identical trainable parameters compared with ViT-S-LLaMA. Concretely, ViT-S-MLP removes the LLM block $\mathbf{F}_{LM}$, and then inserts a GeLU activation (Hendrycks & Gimpel, 2016) and layer normalization (Ba et al., 2016) between $\mathbf{F}_{L}^{1}$ and $\mathbf{F}_{L}^{2}$. It also adopts the identical training procedure as ViT and ViT-LLaMA in Sec. 4.1 for a fair comparison. The results are summarized in Table 6: the ViT-S-MLP has better performance than ViT-S due to its increased capacity, but the improvement is only about half of ViT-S-LLaMA. Therefore, the LLM transformer weights are crucial for the improvement and the observed benefits are not mere consequences of increased model capacity. Investigation with more tasks and baselines are in Sec. B.2.
@@
Fine-tuning.   We further verify whether fine-tuning the language transformer (ViT-S-LLaMA-FT) is better than freezing it. As in Table 6, fine-tuning decreases the performance compared with ViT-S-LLaMA. We analyze this phenomenon by visualizing the loss curves in Fig. F and training under an additional setting of 100 epochs. Although fine-tuning improves performance under short training (100 epochs in Table A), it hurts the accuracy when trained sufficiently: in Fig. F, ViT-S-LLaMA-FT shows lower training loss but relatively larger validation loss, which indicates overfitting. Thus, our observation demonstrates the challenges of training large transformers, and we accordingly freeze the LLM transformers in our design because of its simplicity and effectiveness.
@@ 5.2 Varying LLM Transformer Layers
We discover that different LLM transformers influence visual representation learning significantly within our framework, even though they have identical capacity. Specifically, we use transformer blocks from diverse depths of LLaMA-7B (Touvron et al., 2023) and OPT (Zhang et al., 2022) onto ViT-S. The models are trained in the ablation study setting of 100 epochs. More details are in Sec. C.7. As shown in Fig. 2, the layer type significantly changes the performance. These experiments also validate that our framework is applicable to various LLMs and transformer layers, and highlight the importance of selecting proper transformer layers. We additionally observe that the last LLM layers consistently improve the performance although they might not be optimal.
@@ 6 Information Filtering Hypothesis
This section aims to explain how a pre-trained and frozen LLM transformer benefits visual tasks. Intuitively, our hypothesis can be stated as:
Information filtering hypothesis.   A pre-trained LLM transformer functions as a “filter” that distinguishes the informative tokens and amplifies their contribution for the prediction, in the form of enlarged magnitudes or frequencies in the feature activation.
We first derive this hypothesis in the context of image classification (Sec. 6.1), then provide quantitative investigation (Sec. 6.2), and discuss the observation on other tasks (Sec. 6.3). Due to space limits, we include more details in Sec. A and discuss limitations in Sec. A.4.
@@ 6.1 Qualitative Derivation of Information Filtering Hypothesis
As clearly demonstrated in Fig. 3(a), the token activation better captures the regions of target objects after adding the LLM transformer, especially the magnitudes of $\mathbf{F}_{L}^{2}$ and frequencies of $\mathbf{F}_{LM}^{A}$. Their tendency of segmentation is a surprising discovery, because emergent segmentation is only observed in self-supervised learned (Caron et al., 2021) or specially-designed (Darcet et al., 2024; Shi et al., 2023; Yang et al., 2024) ViTs. More importantly, the activation’s concentration on the target object directly supports our hypothesis as evidence of selecting the informative tokens.
Noisy attention scores.   In contrast to the feature activation, attention scores struggle to capture the relevant visual tokens for prediction. We investigate the attention scores between the $\mathtt{CLS}$ token and visual tokens in the last transformer block, which are the last self-attention block in $\mathbf{F}_{E}$ for ViT and the transformer $\mathbf{F}_{LM}$ for ViT-LLaMA, respectively. Ideal attention scores that distinguish the target object should exhibit object segmentation patterns like DINO (Caron et al., 2021). However, supervised ViT models commonly have noisy attention scores (left part in Fig. 3(b)). Although ViT-LLaMA illustrates the ability of emergent segmentation in a few attention heads, most of the attention scores also suffer from scattering and noisiness. These observations contrast the feature activation and indicate that the benefits of LLM transformers cannot be simply attributed to attention scores, since attention scores fail to reliably contribute correct visual tokens.
Deriving the amplification of informative tokens.   According to our visualization in Fig. 3(a), the frozen LLM transformer distinguishes the informative tokens. Intuitively, such tokens naturally benefit the downstream decoding, but this is straightforward only when the decoder directly takes the visual tokens as input. As a counterexample, ViT utilizes a $\mathtt{CLS}$ token for classification, and the visual tokens output by $\mathbf{F}_{L}^{2}$ is not the input to the decoder and always receives zero gradients during training. To bridge the gap in $\mathtt{CLS}$ token scenarios, the second half of our hypothesis is necessary: the frozen LLM transformer amplifies the contribution of informative tokens.
Concretely, the calculation of the $\mathtt{CLS}$ token is,
@@
where $V$ denotes visual tokens, and $w_{v}$ denotes the weight of visual token $v$. Eqn. 3 describes the process of (1) aggregating visual tokens $z_{L}^{1}[v]$ guided by the attention scores $w_{v}$; and (2) sequentially flowing through the subsequent layers, including the LLM transformer’s attention head $\mathbf{F}_{LM}^{A}$, feedforward network $\mathbf{F}_{LM}^{F}$, and the second linear layer $\mathbf{F}_{L}^{2}$. To concentrate on explaining visual tokens, Eqn. 3 also slightly simplifies self-attention by removing the $\mathtt{CLS}$ token from the right-hand side.
In Eqn. 3, the useful visual tokens are not reliably attributed to the $\mathtt{CLS}$ token, because the attention scores $w_{v}$ are observed to be noisy. Therefore, our objective is to connect the visual tokens in Eqn. 4,
$z_{L}^{2}[v]=\mathbf{F}_{L}^{2}\cdot\mathbf{F}_{LM}^{F}\cdot\mathbf{F}_{LM}^{A}(z_{L}^{1}[v]),\ \text{where}\ v\in V,$ (4)
which are informative (as in Fig. 3(a)), to the final feature representation $z_{L}^{2}[\mathtt{CLS}]$. This pursuit makes us notice that attention scores $w_{v}$ are noisy while token features $z_{L}^{2}[v]$ are informative. Such a contradiction indicates that amplification of the informative tokens is a more plausible explanation, compared with better attention scores. We express the hypothesis below in a formal way, which is a simple change to Eqn. 3:
@@
Eqn. 5 holds equal under the special case of $\mathbf{F}_{L}^{2}\cdot\mathbf{F}_{LM}^{F}\cdot\mathbf{F}_{LM}^{A}$ being linear transformation. This equation explains how the informative tokens in Fig. 3(a) are implicitly supervised in the $\mathtt{CLS}$ token.
As a brief remark, our derivation builds upon two observed pieces of evidence: (1) visual tokens concentrating on informative regions; and (2) noisy attention scores. These lead to the first half of our hypothesis and explain the benefits when visual tokens are direct input to decoders. By further connecting visual tokens to the $\mathtt{CLS}$ token, we propose that the inserted LLM transformer amplifies the effects of informative tokens. A more thorough version of the derivation is in Sec. A.1.
@@ 6.2 Quantitative Evidence
Figure 4: Pseudo-masks from ViT-LLaMA’s features ($\mathbf{F}_{L}^{2}$) have larger mIoU than attention scores and ViT.
@@
Finally, we summarize the mIoU statistics for feature activation and attention scores in Fig. 4. As demonstrated, both ViT-S-LLaMA and ViT-B-LLaMA have better mIoU of pseudo-masks than attention scores. This directly supports our hypothesis that the features $\{z_{L}^{2}[v])|v\in V\}$ contribute more reliably than attention scores $\{w_{v}|v\in V\}$. We additionally notice that the pseudo-masks from ViT-LLaMA generally have larger mIoU compared with ViT, which reflects the benefits of training ViTs with a frozen LLM transformer. The advantage of the feature in the first linear layer $\mathbf{F}_{L}^{1}$ also reveals that training with our framework is beneficial to even earlier stages of features. However, we would like to clarify that the pseudo-masks from either magnitude or frequency activation are intuitive but lossy measures to quantify feature quality, because neural networks can encode information in other formats. Therefore, better measurements to analyze network layers will be meaningful for future work.